# The effect of friction on the dynamics of targeted energy transfer by symmetric vibro-impact dampers

### v1

@@ 1 Introduction
The design of mechanical structures subjected to dynamic loads, such as turbines, aircraft, civil engineering structures and rotating machinery, relies heavily on understanding and controlling vibrations. External forcing or dynamical interactions between components can often induce undesired vibratory behavior potentially leading to malfunctions or total failure of mechanical structures[1, 2]. Given these risks, structural vibration control has become a focal point in ensuring the reliability and longevity of mechanical systems [3, 4]. Whether through modifying the physical properties of the system, e.g. added damping, or through active or passive control [5, 6, 7, 8, 9], all strategies and techniques aim to dissipate unwanted vibrational energy efficiently. A widely studied passive energy control strategy is targeted energy transfer (TET) [10, 9], which directs vibrational energy irreversibly from a primary structure to a nonlinear light attachment known as a nonlinear energy sink (NES). A particularly effective class of NES, especially for high-amplitude oscillations where traditional damping methods fall short, is the vibro-impact nonlinear energy sink (VI-NES)[11, 12]. This absorber mitigates vibrations primarily through impacts, attenuating both stationary and transient responses over a broad frequency range. VI-NES has been investigated in various configurations, including single-sided, double-sided, rotational and inclined setups as well as coupled to cubic nonlinearities [7, 13, 14, 15, 16, 17, 18]. Most prior studies have primarily focused on identifying the most effective operating conditions for TET [19, 20]: a frictionless, symmetric VI-NES exhibiting two symmetric impacts per excitation cycle near primary resonance. This configuration serves as an ideal starting point, as the absence of friction and the symmetry properties of the VI-NES simplify the analysis by isolating the role of impacts on system behavior. Optimal design rules have been derived for these cases [21, 22, 23, 24], providing guidelines for selecting key VI-NES parameters, such as mass ratio, cavity length, and coefficient of restitution, based on the oscillation amplitude and frequency range. However, these models fail to capture the complexities of real-world conditions. In practice, friction plays a significant role in VI-NES behavior, yet previous studies have either not addressed friction in detail or have only considered it for specific setups or only to a limited extent [25, 26, 27, 28, 29, 30]. A systematic analysis of the effects of friction on response regimes, activation thresholds, and the stability of periodic solutions is still lacking. In fact, the presence of friction introduces additional response regimes, which are critical for system design and optimization. Specifically, the interaction between impact-dominated and sliding-dominated behaviors remains insufficiently explored.
This study aims to enhance the predictability of energy dissipation in vibro-impact systems by investigating the dynamics of a linear oscillator (LO) coupled to a symmetric, horizontally oriented VI-NES subjected to dry friction. The coupling is governed by both impact and friction forces, and the study focuses on the influence of these forces on system behavior. Key contributions of this work include the development of a refined analytical framework for the approximation of periodic solutions and their stability, incorporating frictional effects. The study starts by investigating the optimal response regime of two symmetric impacts per period (2 symmetric IPP) under frictional conditions and exploring the dynamics and stability of high-order periodic solutions. It employs complementary approaches, including an extension of the Multiple Scales Method (MSM)[21, 31, 8, 12] to include friction, providing a precise characterization of slow dynamics and the slow invariant manifold (SIM) [32]. Additionally, the impact map approach [33, 34, 26] is introduced to capture fast dynamics, facilitating the exploration of periodic orbits of higher-order. The study also explores various response regimes, including impact-dominated and sliding-dominated behaviors, emphasizing their coexistence and transitions. These contributions collectively provide a deeper understanding of VI-NES dynamics and enable more accurate approximations of frequency response curves (FRCs), offering valuable analytical and numerical tools for the design and optimization of vibro-impact systems in practical applications.
@@
This paper is organized as follows: Section 2 presents the model description, detailing the modifications introduced to incorporate friction as well as the relevant dimensionless parameters. Section 3 extends MSM to frictional VI-NES, deriving the slow dynamics and SIM for symmetric 2IPP solutions. Section 4 introduces the impact map approach, emphasizing its ability to capture fast system dynamics and predict bifurcations and high-order periodic solutions. Section 5 discusses the application of MSM to approximate frequency response curves and identifies the coexistence of distinct response regimes, including impacting and sliding regimes. Section 6 concludes with key findings and future research directions for optimizing VI-NES performance in practical applications.
@@ 2 Model description
The superscripts $(\cdot)^{-}$ and $(\cdot)^{+}$, denote the values at the time instant right before and right after the impact, respectively, and a dot above a variable represents the differentiation with respect to time $t$. The complementarity requirement in (3) guarantees a velocity jump at the moment of impact with a positive impulse preventing the adhesion of the absorber to the side walls of the cavity. Choosing a coefficient of restitution $10\leq r<$ 1, enhances the dissipative character of the VI-NES and adds to the frictional dissipation between impacts a dissipation of energy with every occurring impact.
@@
$\begin{split}&|w|=1:\\ &v^{+}=v^{-}\,,\quad v^{\prime+}=v^{\prime-}\\ &w^{+}=w^{-}\,,\quad w^{\prime+}=-rw^{\prime-}\,,\end{split}$ (10)
@@ 3 Multiple scales method
$\begin{split}&|w_{0}|<1\,,\,\,|D_{0}w_{0}|>\mathcal{O}(\epsilon):\\ &D_{0}^{2}v_{0}+v_{0}=0\,,\\ &D_{0}^{2}w_{0}-D_{0}^{2}v_{0}=-\mu\tilde{g}\operatorname{sign}(D_{0}w_{0})\,,\end{split}$ (15)
@@ 3.2 The slow invariant manifold
[…]
@@ 3.3 Steady-state solutions and flow on the SIM
obtained by using (29) and (30) to replace the sine and cosine terms in (27)-(28), fulfilling the conditions of symmetric two impacts per period (2IPP) motion. Equations (33) describe the phase and amplitude dynamics and are used to determine the existence and stability properties of equilibria on the SIM, corresponding to symmetric periodic motions of the primary and auxiliary mass with 2IPP. Setting the right-hand side of (33) to zero, as to find equilibria, reduces the problem after elimination of $\gamma$ to a single quadratic algebraic equation of the form
@@
Solving (34) for positive real roots delivers a closed-form analytical expression of the steady-state squared amplitude $\tilde{C}$ as a function of the velocity $B$. The obtained expression describes another manifold in the slow parameter configuration space, that unlike the SIM from (31), depends on the external excitation through the parameters $G$ and $\sigma$, as well as the damping constant $\lambda$. Thus, steady-state solutions with 1:1 resonance and 2 symmetric impacts per period can be identified graphically as the intersection of the two manifolds in the $(B,\,\tilde{C})$-plane. A graphical illustration is given in Figure (5) for different excitation frequencies and the same level of forcing. One should note that only points on the purple curves which agree with the SIM have a physical meaning and represent the equilibria of (33). The stability of the equilibria on the SIM is determined by examining the eigenvalues of the Jacobian matrix $\mathcal{J}$ of the right-hand sides of (33), evaluated at the equilibrium solution, after expressing the velocity $B$ as
@@
$\begin{split}\tilde{\alpha}_{c}&=\lambda+\frac{4}{\pi}\frac{R}{\frac{\pi^{2}}{4}+R^{2}}\,,\\ \tilde{\alpha}_{\gamma}&=\frac{2}{\frac{\pi^{2}}{4}+R^{2}}+2\sigma\,,\\ \tilde{\beta}_{c}&=\frac{1}{\sqrt{\frac{\pi^{2}}{4}+R^{2}}}\left(\frac{8}{\pi}RB_{\min}+\frac{16+8\pi^{2}-\pi^{4}}{2\pi^{3}}h\right)\,,\\ \tilde{\beta}_{\gamma}&=\frac{1}{\sqrt{\frac{\pi^{2}}{4}+R^{2}}}\left(4B_{\min}-\frac{4}{\pi}-\frac{16-\pi^{4}}{\pi^{4}}Rh\right)\,,\\ \tilde{K}_{c}&=-\frac{4}{\pi}R\left(B_{\min}^{2}-\frac{\tilde{C}_{\min}}{\frac{\pi^{2}}{4}+R^{2}}\right)-\frac{16+8\pi^{2}-\pi^{4}}{2\pi^{3}}hB_{\min}+\frac{16-\pi^{4}}{\pi^{4}}h\,,\\ \tilde{K}_{\gamma}&=\frac{4}{\pi}B_{\min}\left(1-\frac{\pi}{2}B_{\min}\right)+\frac{2}{\frac{\pi^{2}}{4}+R^{2}}\tilde{C}_{\min}+\frac{16-\pi^{4}}{\pi^{4}}h\left(RB_{\min}+h\right)\,,\end{split}$ (39)
@@
Once on the SIM, the system’s slow dynamics dictate the stability of steady-state periodic solutions, such as symmetric 2IPP motion. A stable equilibrium on the SIM ensures that nearby trajectories remain close, enabling sustained periodic energy transfer between the primary system and the NES. However, if the equilibrium is unstable, the slow flow diverges from it, potentially leading to chaotic or transient responses. However, these equations do not provide any information about the attractiveness of the SIM itself. If the SIM is attractive, trajectories in its vicinity will asymptotically converge to it, even if the system starts off the SIM. Notably, it is possible for the slow dynamics to possess an unstable equilibrium located on the stable branch of the SIM. The behavior of the system in such cases will be discussed in Section 5. The upcoming section will carry out the stability and bifurcation analysis for the SIM, identifying the set of all possible solutions of the motion under consideration and determining the corresponding attractive branches.
@@ 4.2 The elementary mappings
The obtained elementary mappings can be iterated to generate various types of motion with alternating impacts, specifically the $2$IPP motion examined in this work. This motion can be described by two implicit impact maps, $\mathbf{G}^{+}$ or $\mathbf{G}^{-}$, constructed from the derived elementary mappings from Table 1 as follows:
@@
It is important to note that the derived elementary mappings in this subsection describe alternating impacts. However, the auxiliary mass may also impact the same sidewall twice without reaching the opposite wall. These scenarios, along with the corresponding elementary mappings, denoted by $\mathbf{G}^{\pm}_{\text{LL}}$ and $\mathbf{G}^{\pm}_{\text{RR}}$, were derived and extensively analyzed in [36] and therefore fall beyond the scope of this paper, which focuses exclusively on the $2$IPP motion.
@@ 4.3 Periodic two impacts per cycle motion: $\mathcal{P}_{2}^{1}$-orbits
The definition of an $l$-periodic orbit of order $k$, as introduced in [33], characterizes an orbit of the impact map $\mathbf{G}^{\pm}$ with a period of $\frac{2\pi}{\Omega}k$ and $l$ impacts occurring within one period of time ($1$:$k$ internal resonance with $l$ impacts). The $\mathcal{P}_{2}^{1}$-orbits, with two impacts per cycle in the continuous time representation, correspond to period-1 fixed points of the impact maps $\mathbf{G}^{\pm}$ from equation (52). Such fixed points, in the following denoted by
@@
$\mathbf{G}^{+}\left(\mathbf{x}^{+*}_{n},\,\mathbf{x}^{+*}_{n+2}\right)=\mathbf{0}\Rightarrow\,\left\{\begin{array}[]{ll}C\cos\Psi^{*}&=R(B^{+*}-\tilde{h})+\tilde{h}\\ C\sin\Psi^{*}&=1-\frac{\pi}{2}\left(B^{+*}-\tilde{h}\right)\end{array}\right.$ (59)
$\mathbf{G}^{-}\left(\mathbf{x}^{-*}_{n},\,\mathbf{x}^{-*}_{n+2}\right)=\mathbf{0}\Rightarrow\,\left\{\begin{array}[]{ll}C\cos\Psi^{*}&=R(B^{-*}+\tilde{h})+\tilde{h}\\ C\sin\Psi^{*}&=1-\frac{\pi}{2}\left(B^{-*}+\tilde{h}\right)\end{array}\right.$ (62)
with $\tilde{h}=h(1+\epsilon)$ and the parameters $R$ and $h$ are defined according to (31). Using the trigonometric identity yields the following polynomial expressions that relate the post- and pre-impact velocities to the squared value of the external excitation amplitude during the symmetric 2 IPP motion
@@ 4.4 Linear stability and bifurcation analysis
$\begin{split}\mathbf{0}&=\mathbf{G}\left(\mathbf{x}_{n},\,\mathbf{x}_{n+2}\right)=\begin{pmatrix}\mathbf{G}_{\text{LR}}\left(\mathbf{x}_{n},\,\mathbf{x}_{n+1}\right)\\ \mathbf{G}_{\text{RL}}\left(\mathbf{x}_{n+1},\,\mathbf{x}_{n+2}\right)\\ \end{pmatrix}\\ &=\begin{pmatrix}\mathbf{G}_{\text{LR}}\left(\mathbf{x}^{*}_{n},\,\mathbf{x}^{*}_{n+1}\right)+\frac{\partial\mathbf{G}_{\text{LR}}}{\partial\mathbf{x}_{n}}\left(\mathbf{x}^{*}_{n},\,\mathbf{x}^{*}_{n+1}\right)\Delta\mathbf{x}_{n}\cdots\\ \cdots+\frac{\partial\mathbf{G}_{\text{LR}}}{\partial\mathbf{x}_{n+1}}\left(\mathbf{x}^{*}_{n},\,\mathbf{x}^{*}_{n+1}\right)\Delta\mathbf{x}_{n+1}+\cdots\\ \\ \mathbf{G}_{\text{RL}}\left(\mathbf{x}^{*}_{n+1},\,\mathbf{x}^{*}_{n+2}\right)+\frac{\partial\mathbf{G}_{\text{RL}}}{\partial\mathbf{x}_{n+1}}\left(\mathbf{x}^{*}_{n+1},\,\mathbf{x}^{*}_{n+2}\right)\Delta\mathbf{x}_{n+1}\cdots\\ \cdots+\frac{\partial\mathbf{G}_{\text{RL}}}{\partial\mathbf{x}_{n+2}}\left(\mathbf{x}^{*}_{n+1},\,\mathbf{x}^{*}_{n+2}\right)\Delta\mathbf{x}_{n+2}+\cdots\end{pmatrix}\,.\end{split}$ (68)
@@
$\begin{split}a&=\pi\tilde{k}-\frac{1-R}{1+R}\,,\\ b&=\tilde{k}\left(2\tilde{h}-(1-R)(B^{\pm}\mp\tilde{h})\right)-\frac{1-R}{1+R}\frac{2}{\pi}\tilde{h}\,,\\ c&=\frac{-\pi}{\left(1+R\right)\left(B^{\pm}\mp\tilde{h}\right)}\,,\\ d&=\frac{1-R}{1+R}-\frac{2\tilde{h}}{\left(1+R\right)\left(B^{\pm}\mp\tilde{h}\right)}\,,\\ \text{with}\quad\tilde{k}&=\frac{2}{(1+R)^{2}}\left(\frac{\pi}{2}-\frac{1}{B^{\pm}\mp\tilde{h}}\right)-\frac{(1-R)\frac{2}{\pi}\tilde{h}}{\left(1+R\right)^{2}\left(B^{\pm}\mp\tilde{h}\right)}\,,\end{split}$ (74)
@@
[…]
@@
If $\sin(\phi)=0$, then both eigenvalues are situated on the real axis. In this case, instability can arise only if one of the eigenvalues crosses the value $+1$ or $-1$. Taking this into consideration, along with stability requirement $\rho<1$, equation (77) delivers the condition
@@
$\begin{split}B^{\pm}_{\min}<B^{\pm}<B^{\pm}_{\max}&\\ \text{with}\quad B^{\pm}_{\min}=\frac{2\pi-4R\tilde{h}}{\pi^{2}+4R^{2}}\pm\tilde{h}\quad\text{and}\quad B^{\pm}_{\max}&=\frac{2\pi-4R\tilde{h}}{\pi^{2}-4}\pm\tilde{h}\,.\end{split}$ (84)
@@
The expression for the activation threshold $\tilde{C}_{\min}$ reveals its dependence not only on the restitution coefficient $r$ (through $R$), the friction coefficient $\mu$ and the cavity length $b$ (through $\tilde{h}$), as in the MSM approach, but also on the mass ratio $\epsilon$. Furthermore, local asymptotic stability of symmetric $\mathcal{P}_{2}^{1}$- orbits is guaranteed only for excitation amplitudes within the range $(\tilde{C}_{\min},\tilde{C}_{\max})$, where the average absolute velocity of the absorber satisfies the inequality $B_{\min}<B<B_{\max}$, with the inequality bounds from (85). If the amplitude $\tilde{C}$, considered here as the changing bifurcation parameter, goes beyond these bounds, the fixed points lose their stability, potentially leading to the emergence of new solutions. The nature of the new motion depends on the type of bifurcation that occurs, which is determined by how the eigenvalues leave the unit circle. According to (75), the eigenvalues cannot leave the unit circle as a complex conjugate pair, since their product has been shown to satisfy $\lambda_{1}\lambda_{2}=r^{4}<1$. Therefore, bifurcations occur only for real-valued eigenvalues. If an eigenvalue of the linearized map evaluated at the fixed point moves through $+1$, then one of the following bifurcations can occur: first, a saddle-node bifurcation, characterized by the simultaneous creation or destruction of stable and unstable solution branches. The corresponding bifurcation point represents a boundary for the existence of solutions in the parameter space. The second type of bifurcation is a symmetry-breaking bifurcation, characterized by the creation of asymmetric solutions from symmetric ones. When an eigenvalue is at $-1$, only a period-doubling bifurcation occurs. In this case, the new steady-state motion has twice the period of the one it emerges from. In this context, solving $p_{\mathbf{A}}(\lambda=+1)=0$ results in the same values presented in (84) and subsequently (87). This means that the bifurcations that occur on the SIM are either a saddle-node, or a symmetry-breaking bifurcation. However, period-doubling bifurcations do not occur on the SIM, since $p_{\mathbf{A}}(\lambda=-1)=0$ has no solutions.
Figure 9: Stability region of the $\mathcal{P}_{2}^{1}$-orbits for $r=0.76$, $\mu=0.5$ and $\tilde{g}=0.211$. The black lines result from $\mathbf{G}^{+}\left(\mathbf{x}_{n},\,\mathbf{x}_{n+2}\right)=\mathbf{0}$, $\mathbf{G}^{-}\left(\mathbf{x}_{n},\,\mathbf{x}_{n+2}\right)=\mathbf{0}$ and (65). The green lines correspond to the stable branches.

### v2

@@ 1 Introduction
This paper presents a unified analytical and numerical framework for characterizing the dynamics of a symmetric vibro-impact nonlinear energy sink with dry friction. By extending the multiple scales method to include friction and combining it with an impact-map approach, closed-form expressions for performance metrics are derived linking tuning parameters, including friction, to activation, dissipation, and stability within the two-impacts-per-period regime. The analysis reveals coexisting impact- and sliding-dominated responses near the optimal operating range and clarifies the amplitude-dependent role of friction in targeted energy transfer. The design of mechanical structures subjected to dynamic loads, such as turbines, aircraft, civil engineering structures, and rotating machinery, relies heavily on understanding and controlling vibrations. External forcing or dynamical interactions between components can often induce undesired vibratory behavior potentially leading to malfunctions or total failure of mechanical structures [1, 2]. Given these risks, structural vibration control has become a focal point in ensuring the reliability and longevity of mechanical systems [3, 4]. Whether through modifying the physical properties of the system, e.g. added damping, or through active or passive control [5, 6, 7, 8, 9], all strategies and techniques aim to dissipate unwanted vibrational energy efficiently. A widely studied passive energy control strategy is targeted energy transfer (TET) [10, 9], which directs vibrational energy irreversibly from a primary structure to a nonlinear light attachment known as a nonlinear energy sink (NES). NES configurations have been continuously optimized to improve energy-transfer efficiency and broaden their effective operating range. Among these, bistable NES (BNES) designs using negative-stiffness elements have demonstrated high efficiency in vibration absorption [11, 12]. More recent studies combining bistability with vibro-impact mechanisms have shown enhanced performance, improved robustness, and extended operational bandwidth [13, 14]. The use of BNES, however, is restricted to devices in which negative stiffness effects can be realized.
A particularly effective class of NES, especially for high-amplitude oscillations where traditional damping methods fall short, is the vibro-impact nonlinear energy sink (VI-NES) [15, 16]. This absorber mitigates vibrations primarily through impacts, attenuating both stationary and transient responses over a broad frequency range.
A typical VI-NES consists of a lightweight auxiliary mass that moves freely within a cavity and collides with its boundaries. This concept has been explored in various configurations [15, 17], including single- and double-sided designs [18, 19, 20], inclined [21, 22, 23], and rotational variants [24, 25], as well as setups incorporating additional cubic nonlinearities [7, 26]. Due to their consistent performance across diverse excitation scenarios [27, 28], VI-NES have also been adapted to more complex setups and practical applications, such as seismic mitigation [1, 18], vibration suppression under traveling loads [29], and vibration-based energy harvesting [16, 30]. Most prior studies have primarily focused on identifying the most effective operating conditions for TET [31, 32], by analyzing a frictionless, symmetric VI-NES exhibiting two symmetric impacts per excitation cycle near primary resonance. This idealized configuration serves as a logical starting point, as the absence of friction and the symmetry properties of the VI-NES simplify the analysis by isolating the role of impacts on system behavior and allowing a clear analytical characterization of the most effective response regime that marks the onset of TET and the activation of the VI-NES. Optimal design rules have been derived for these cases [33, 34, 35, 36], providing guidelines for selecting key VI-NES parameters, such as mass ratio, cavity length, and coefficient of restitution, based on the oscillation amplitude and frequency range. However, such idealized models fail to capture the complexities of real-world behavior, where friction plays a crucial role in shaping the VI-NES response. Previous studies that included friction have only examined specific configurations or investigated the frictional effects to a limited extent, either experimentally or through numerical parametric analyses [37, 38, 21, 39, 40, 41]. Consequently, a systematic analytical framework that quantifies the influence of friction on response regimes, its interaction with other tuning parameters, and its impact on overall VI-NES performance is still lacking. In practice, the presence of friction introduces additional response regimes, which are critical for system design and optimization. Specifically, the interaction between impact-dominated and sliding-dominated behaviors remains insufficiently explored, despite its significant influence on energy-transfer efficiency and absorber robustness.
This study aims to enhance the predictability of energy dissipation in vibro-impact systems by analyzing the dynamics of a linear oscillator (LO) coupled to a symmetric, horizontally oriented VI-NES subjected to dry friction. The coupling is governed by both impact and friction forces, and the work focuses on their combined influence on system behavior. Building upon the analytical framework developed in our previous study of the symmetric frictionless VI-NES [33], the present work extends the Multiple Scales Method (MSM) [33, 42, 8, 16, 43] to explicitly account for frictional effects. This extension enables the derivation of closed-form expressions for design-relevant performance metrics, thereby providing a new analytical approach for evaluating VI-NES performance under realistic conditions. To complement the MSM, an impact-map approach [44, 45, 38] is employed to capture the fast dynamics, determine the stability boundaries of the two-impacts-per-period ($2$-IPP) regime, and fully characterize the optimal operating range of the frictional VI-NES. The global behavior, including the emergence of higher-order impact sequences and the associated bifurcations, was investigated in a recent study [46] for the general case of an asymmetric VI-NES with dry friction. Together, these contributions establish a comprehensive analytical and numerical framework that deepens the understanding of VI-NES dynamics, clarifies the interplay between friction, impacts, and tuning parameters, and provides practical tools for the design and optimization of vibro-impact nonlinear energy sinks.
@@
This paper is organized as follows: Section 2 presents the model description, detailing the modifications introduced to incorporate friction as well as the relevant dimensionless parameters. Section 3 extends MSM to frictional VI-NES, deriving the slow dynamics and SIM for symmetric $2$-IPP solutions. Section 4 introduces the impact-map approach, emphasizing its ability to capture fast system dynamics and predict the stability and bifurcation properties of the studied periodic solutions. Section 5 synthesizes the results from the MSM and impact-map analyses to define design-relevant metrics that quantify the performance and efficiency of the VI-NES, followed by a discussion of the coexistence between impacting and purely sliding response regimes and the approximation of the corresponding frequency response curves based on the MSM formulation. Section 6 concludes with key findings and future research directions for optimizing VI-NES performance in practical applications.
@@ 2 Model description
The superscripts $(\cdot)^{-}$ and $(\cdot)^{+}$, denote the values at the time instant right before and right after the impact, respectively, and a dot above a variable represents the differentiation with respect to time $t$. The complementarity requirement in (3) guarantees a velocity jump at the moment of impact with a positive impulse preventing the adhesion of the absorber to the side walls of the cavity. Choosing a coefficient of restitution $0\leq r<1$, enhances the dissipative character of the VI-NES and adds to the frictional dissipation between impacts a dissipation of energy with every occurring impact.
@@
[…]
@@ 3 Multiple scales method
[…]
@@ 3.2 The slow invariant manifold
$\textbf{Contact condition:}\;\left|w_{0}(\tau^{c}_{0,k})\right|=1$ $\Rightarrow C\cos\left(\eta-\theta\right)$ $=1-\frac{\pi}{2}B\,$, (29)
$\textbf{Impact equation:}\;D_{0}^{+}w_{0}=-rD_{0}^{-}w_{0}$ $\Rightarrow C\sin\left(\eta-\theta\right)$ $=\frac{1-r}{1+r}B+\frac{\pi}{2}\mu\tilde{g}\,$. (30)
@@ 3.3 Steady-state solutions and flow on the SIM
obtained by using (29) and (30) to replace the sine and cosine terms in (27)-(28), fulfilling the conditions of symmetric two impacts per period ($2$-IPP) motion. Equations (33) describe the phase and amplitude dynamics and are used to determine the existence and stability properties of equilibria on the SIM, corresponding to symmetric periodic motions of the primary and auxiliary mass with $2$-IPP. Setting the right-hand side of (33) to zero, as to find equilibria, reduces the problem after elimination of $\gamma$ to a single quadratic algebraic equation of the form
@@
[…]
@@
[…]
@@
Once on the SIM, the system’s slow dynamics dictate the stability of steady-state periodic solutions, such as symmetric $2$-IPP motion. A stable equilibrium on the SIM ensures that nearby trajectories remain close, enabling sustained periodic energy transfer between the primary system and the NES. However, if the equilibrium is unstable, the slow flow diverges from it, potentially leading to chaotic or transient responses. However, these equations do not provide any information about the attractiveness of the SIM itself. If the SIM is attractive, trajectories in its vicinity will asymptotically converge to it, even if the system starts off the SIM. Notably, it is possible for the slow dynamics to possess an unstable equilibrium located on the stable branch of the SIM. The behavior of the system in such cases will be discussed in Section 5. The upcoming section will carry out the stability and bifurcation analysis for the SIM, identifying the set of all possible solutions of the motion under consideration and determining the corresponding attractive branches.
@@ 4.2 The elementary mappings
The obtained elementary mappings can be iterated to generate various types of motion with alternating impacts, specifically the $2$-IPP motion examined in this work. This motion can be described by two implicit impact maps, $\mathbf{G}^{+}$ or $\mathbf{G}^{-}$, constructed from the derived elementary mappings from Table 1 as follows:
@@
It is important to note that the derived elementary mappings in this subsection describe alternating impacts. However, the auxiliary mass may also impact the same sidewall twice without reaching the opposite wall. These scenarios, along with the corresponding elementary mappings, denoted by $\mathbf{G}^{\pm}_{\text{LL}}$ and $\mathbf{G}^{\pm}_{\text{RR}}$, were derived and extensively analyzed in [46] and therefore fall beyond the scope of this paper, which focuses exclusively on the $2$-IPP motion.
@@ 4.3 Periodic two impacts per cycle motion: $\mathcal{P}_{2}^{1}$-orbits
The definition of an $l$-periodic orbit of order $k$, as introduced in [44], characterizes an orbit of the impact map $\mathbf{G}^{\pm}$ with a period of $\frac{2\pi}{\Omega}k$ and $l$ impacts occurring within one period of time (11:$k$ internal resonance with $l$ impacts). The $\mathcal{P}_{2}^{1}$-orbits, with two impacts per cycle in the continuous time representation, correspond to period-1 fixed points of the impact maps $\mathbf{G}^{\pm}$ from equation (52). Such fixed points, in the following denoted by
@@
[…]
with $\tilde{h}=h(1+\epsilon)$ and the parameters $R$ and $h$ are defined according to (31). Using the trigonometric identity yields the following polynomial expressions that relate the post- and pre-impact velocities to the squared value of the external excitation amplitude during the symmetric $2$-IPP motion
@@ 4.4 Linear stability and bifurcation analysis
[…]
@@
[…]
@@
$\textbf{Case 1:}\quad\text{Tr}(\mathbf{A})=2\rho\cos(\phi)\,$, (79)
$\textbf{Case 2:}\quad\sin(\phi)=0\,.$ (80)
@@
If $\sin(\phi)=0$, then both eigenvalues are situated on the real axis. In this case, instability can arise only if one of the eigenvalues crosses the value +1+1 or 1-1. Taking this into consideration, along with stability requirement $\rho<1$, equation (77) delivers the condition
@@
[…]
@@
The expression for the activation threshold $\tilde{C}_{\min}$ reveals its dependence not only on the restitution coefficient $r$ (through $R$), the friction coefficient $\mu$ and the cavity length $b$ (through $\tilde{h}$), as in the MSM approach, but also on the mass ratio $\epsilon$. Furthermore, local asymptotic stability of symmetric $\mathcal{P}_{2}^{1}$- orbits is guaranteed only for excitation amplitudes within the range $(\tilde{C}_{\min},\tilde{C}_{\max})$, where the average absolute velocity of the absorber satisfies the inequality $B_{\min}<B<B_{\max}$, with the inequality bounds from (85). If the amplitude $\tilde{C}$, considered here as the changing bifurcation parameter, goes beyond these bounds, the fixed points lose their stability, potentially leading to the emergence of new solutions. The nature of the new motion depends on the type of bifurcation that occurs, which is determined by how the eigenvalues leave the unit circle. According to (75), the eigenvalues cannot leave the unit circle as a complex conjugate pair, since their product has been shown to satisfy $\lambda_{1}\lambda_{2}=r^{4}<1$. Therefore, bifurcations occur only for real-valued eigenvalues. If an eigenvalue of the linearized map evaluated at the fixed point moves through +1+1, then one of the following bifurcations can occur: first, a saddle-node bifurcation, characterized by the simultaneous creation or destruction of stable and unstable solution branches. The corresponding bifurcation point represents a boundary for the existence of solutions in the parameter space. The second type of bifurcation is a symmetry-breaking bifurcation, characterized by the creation of asymmetric solutions from symmetric ones. When an eigenvalue is at 1-1, only a period-doubling bifurcation occurs. In this case, the new steady-state motion has twice the period of the one it emerges from. In this context, solving $p_{\mathbf{A}}(\lambda=+1)=0$ results in the same values presented in (84) and subsequently (87). This means that the bifurcations that occur on the SIM are either a saddle-node, or a symmetry-breaking bifurcation. However, period-doubling bifurcations do not occur on the SIM, since $p_{\mathbf{A}}(\lambda=-1)=0$ has no solutions.
Figure 9: Stability region of the $\mathcal{P}_{2}^{1}$-orbits for $r=0.76$, $\mu=0.5$ and $\tilde{g}=0.211$. The black lines result from $\mathbf{G}^{\pm}\left(\mathbf{x}_{n},\,\mathbf{x}_{n+2}\right)=\mathbf{0}$ and (65). The green lines correspond to the stable branches.